# AfriDoc-MT: Document-level MT Corpus for African Languages

### v1

@@ C.2 Document-level evaluation
For document-level evaluation, we split the documents into chunks of 10 sentences and translate these chunks using the different models. In Tables 17 and 18 we provide the full results on the merged pseudo-documents using d-CHRF and d-BLEU. And below are some other relevant points from the results. It is important to note that we also trained and evaluated NLLB-200 for pseudo-document translation; however, due to its 512-token maximum sequence length, it is not competitive. Nevertheless, the results still show the influence of fine-tuning. Below are other findings.
@@
Finetuning data has impact of translation quality.
Our results show that both LLama3.1 and LLaMAX3 models, when finetuned on sentences, performed significantly worse on pseudo-document evaluations compared to the same models finetuned on pseudo-documents for both domains. All these models were trained using a similar setup, with the primary difference being the data used for finetuning.
C.3 GPT as a proxy for human evaluation
In Tables 8 and 12 We present the GPT-4o evaluation results for five models, including the GPT-4o translation outputs for both domains, evaluating translations between English and three African languages: Hausa, Swahili and Zulu due to resource constraint. The results show that GPT-4o achieves the best overall performance, demonstrating high fluency, fewer content errors, and fewer lexical and grammatical errors, which can be attributed to self-bias. However, our findings indicate that, overall, document-level translation output (pseudo-documents) are more fluent compared to sentence-level translations. Similarly, document-level translations have fewer lexical and grammatical errors, although content errors are not specifically consistent.
@@ Appendix D More discussion and analysis
A manual inspection of the outputs of GPT-4o when used as a proxy for human evaluation in evaluating the translated documents obtained for selected LLMs and language pairs reveals several issues. These include word and phrase repetition, off-target translations, incorrect translations, and possible omissions. These issues, which are also identified in the literature as common in document-level translation with LLMs Wu et al. (2024); Wang et al. (2024b). To gain a better understanding of some of these issues, we conducted an analysis by computing statistics on the model’s pseudo-document outputs and posing specific questions.
We analyzed the translation outputs comparing them to their corresponding reference translation to check wether they are, empty or if the models under-generate. Our analysis shows that all models rarely generate empty translations (refer to Figure 8). However, GPT-3.5 and GPT-4o exhibit a slight tendency to generate empty translations when translating into Yorùbá and Zulu for both domains, though this occurs rarely with a frequency of 10%. For output length, translations with lengths less than 70% of the ground truth were considered under-generated. Figures 9 and 10 shows Aya-101 under-generates more than 3% of the pseudo-documents when translating into African languages. In the other direction, LLaMAX-Alpaca shows atleast 15% under-generation into English, while other models have considerable amount of under-generation as well.
When we compare our SFT models trained on sentences and pseudo-documents with k=10, our result in Figure 13 shows that the models trained on sentences under-generates when used for long document translation.
Do LLMs generate translations in the correct target languages?
We evaluate whether these models understand the task by generating outputs in the target languages using the OpenLID (Burchell et al., 2023) language identification model. Our results show that these models rarely generate outputs in the wrong language when translating to English. However, when translating to African languages, there is a higher likelihood of incorrect language translations, particularly with open models (Figure 11).
What is the effect of document length on translation quality?
We compare the average d-CHRF scores obtained by selected models, including GPT-3.5/4 and LLama3.1-SFTk where 1,5,10. The evaluation was conducted across all pseudo-document lengths: 1, 5, 10, 25, and the full length. Figure 3 shows that for translations into African languages, d-CHRF scores decrease as document length increases. A similar trend is observed for the reverse translation, except for GPT-4o, which shows an increasing trend.
@@
We focus on the sentence-level task and translated across all 30 directions for which the model was trained, evaluating both NLLB-200 (1.3B) and its fine-tuned version using d-CHRF. Figures 14 and 15 show performance improvements after supervised fine-tuning of NLLB-200 for both domains. The results shows that translating into Yorùbá, which is the direction with the lowest d-CHRF score from English among all the languages, benefited the most. One major factor contributing to this is the presence of diacritics.

### v2

@@ C.2 Document-level evaluation
For document-level evaluation, we split the documents into chunks of 10 sentences and translate these chunks using the different models. In Tables˜19 and 20 we provide the full results on the merged pseudo-documents using d-chrF and d-BLEU. And below are some other relevant points from the results. It is important to note that we also trained and evaluated NLLB-200 for pseudo-document translation. However, due to its 512-token maximum sequence length, it is not competitive. Nevertheless, the results still show the influence of fine-tuning. Below are other findings.
@@
Fine-tuning data has an impact on translation quality.
Our results show that both LLama3.1 and LLaMAX3 models, when fine-tuned on sentences, performed significantly worse on pseudo-document evaluations compared to the same models fine-tuned on pseudo-documents for both domains. All these models were trained using a similar setup, with the primary difference being the data used for fine-tuning.
Language-specific performance trends
Overall, no clear trend is observed in MT performance across language family classes. However, Amharic (a non-Latin script language) and Yorùbá (a heavily diacriticitized language) result in the lowest chrF scores, while Swahili—the most widely spoken indigenous African language—performs best.
C.3 Findings from GPT-4o as a judge
In Tables˜21 and 22 we present the average GPT-4o evaluation results for four models. When translating into African languages, there is no clear pattern: for example, GPT-3.5, despite having the lowest fluency score, also had the fewest content, lexical, and grammatical errors, which is counterintuitive. In contrast, when translating into English, the pattern is clear and consistent: translations of pseudo-documents show better fluency and fewer errors overall. These results suggest that using GPT-4o as a translation judge is not yet well-suited for low-resource languages.
| Model | Setup | amh | hau | swh | yor | zul |
| GPT-3.5 | Sent | 18.3 | 42.1 | 64.1 | 12.6 | 8.9 |
| Doc10 | 4.8 | 32.3 | 58.5 | 6.9 | 17.5 | |
| LLaMAX3-SFT1 | Sent | 58.1 | 86.0 | 62.1 | 66.1 | 52.0 |
| Doc10 | 19.0 | 54.5 | 38.7 | 40.1 | 21.6 | |
| LLaMAX3-SFT10 | Doc10 | 54.3 | 83.7 | 61.7 | 62.3 | 37.1 |
C.4 Findings from human evaluation
We were able to obtain DA scores from three annotators for all the languages. For each language, we calculated inter-annotator agreement using Krippendorff’s alpha over 30 document instances. We obtained scores of 0.46, 0.57, 0.40, and 0.81, and 0.54 for Amharic, Hausa, Swahili, Yorùbá, and Zulu respectively. These are relatively low scores, except for Yorùbá. We present the average DA scores in Tables˜10 and 14 for the health and tech domains, respectively. The results show that annotators rate documents translated at the sentence-level as higher quality than those translated at the pseudo-document level. Additionally, GPT-3.5 received the lowest ratings among the three models. LLaMAX3-SFT1, a model trained on sentence-level data, was rated the best across all languages when evaluated on sentences. However, when evaluated on pseudo-documents, its performance was rated lower than that of LLaMAX3-SFT10. These findings are consistent with the d-chrF scores for the models, but they do not align with the evaluations from GPT-4o as a judge.
@@ Appendix D More discussion and analysis
We focus on the sentence-level task and translated across all 30 directions for which the model was trained, evaluating both NLLB-200 (1.3B) and its fine-tuned version using d-chrF. Figures˜13 and 14 show performance improvements after supervised fine-tuning of NLLB-200 for both domains. The results shows that translating into Yorùbá, which is the direction with the lowest d-chrF score from English among all the languages, benefited the most. One major factor contributing to this is the presence of diacritics. Furthermore, looking at their actual performances and not just the differences, our results show that translations into Swahili and English—both relatively high-resource languages—yield higher BLEU and chrF scores (see Figures 11 and 12), even after supervised finetuning. Hence, there is much to be done to improve translation performance between low-resource language pairs.
@@
| Model | Setup | eng X | X eng | ||||||||
| d-CHRF | Fluency | CE | LE | GE | d-CHRF | Fluency | CE | LE | GE | ||
| Amharic | |||||||||||
| Aya-101 | Sent | 36.6 | 2.8 | 9.0 | 3.8 | 2.5 | 64.6 | 3.0 | 15.8 | 9.6 | 5.1 |
| Doc10 | 28.7 | 2.5 | 8.9 | 2.9 | 2.1 | 61.6 | 3.6 | 12.6 | 8.5 | 3.8 | |
| GPT-3.5 | Sent | 20.4 | 1.1 | 7.2 | 10.8 | 3.1 | 48.3 | 3.0 | 9.5 | 3.7 | 2.4 |
| Doc10 | 11.6 | 1.0 | 3.3 | 1.8 | 1.6 | 41.6 | 4.2 | 6.6 | 2.0 | 1.3 | |
| LLaMAX3-SFT1 | Sent | 46.8 | 3.5 | 10.0 | 2.9 | 2.0 | 66.6 | 3.5 | 11.5 | 6.0 | 2.9 |
| Doc10 | 24.1 | 1.7 | 10.4 | 1.9 | 1.8 | 22.6 | 3.0 | 9.0 | 2.5 | 1.6 | |
| LLaMAX3-SFT10 | Doc10 | 37.8 | 2.6 | 7.3 | 1.8 | 1.6 | 64.0 | 4.1 | 8.7 | 4.6 | 2.7 |
| Hausa | |||||||||||
| Aya-101 | Sent | 56.4 | 2.9 | 9.1 | 4.1 | 3.7 | 61.5 | 2.9 | 17.1 | 10.2 | 5.6 |
| Doc10 | 48.5 | 2.9 | 8.3 | 2.3 | 1.9 | 62.3 | 3.2 | 14.6 | 8.4 | 4.5 | |
| GPT-3.5 | Sent | 44.3 | 1.4 | 11.3 | 5.2 | 5.8 | 52.4 | 2.5 | 13.2 | 6.8 | 3.7 |
| Doc10 | 23.1 | 1.1 | 7.4 | 2.9 | 2.8 | 52.7 | 4.1 | 9.2 | 4.1 | 2.3 | |
| LLaMAX3-SFT1 | Sent | 62.5 | 3.2 | 9.5 | 4.0 | 3.5 | 58.9 | 2.9 | 9.6 | 4.1 | 3.0 |
| Doc10 | 29.0 | 2.3 | 8.8 | 2.3 | 1.9 | 22.9 | 2.6 | 9.0 | 3.0 | 1.9 | |
| LLaMAX3-SFT10 | Doc10 | 51.9 | 3.9 | 14.6 | 2.2 | 1.9 | 66.7 | 4.2 | 9.6 | 4.5 | 2.6 |
| Swahili | |||||||||||
| Aya-101 | Sent | 44.7 | 1.3 | 9.5 | 5.0 | 3.4 | 70.8 | 3.3 | 17.5 | 9.9 | 4.8 |
| Doc10 | 34.7 | 1.8 | 8.2 | 5.0 | 3.0 | 71.2 | 3.8 | 14.1 | 9.8 | 4.0 | |
| GPT-3.5 | Sent | 76.7 | 4.9 | 4.2 | 1.8 | 1.2 | 75.0 | 3.6 | 12.4 | 7.6 | 3.7 |
| Doc10 | 76.1 | 4.9 | 3.9 | 0.8 | 0.6 | 77.7 | 4.7 | 7.1 | 4.5 | 2.0 | |
| LLaMAX3-SFT1 | Sent | 73.1 | 3.6 | 12.3 | 4.8 | 3.4 | 73.1 | 3.9 | 11.8 | 8.4 | 2.7 |
| Doc10 | 42.2 | 3.2 | 9.5 | 3.6 | 2.5 | 33.1 | 3.1 | 8.9 | 3.2 | 1.9 | |
| LLaMAX3-SFT10 | Doc10 | 74.4 | 4.4 | 10.5 | 3.6 | 2.2 | 77.8 | 4.6 | 8.9 | 6.2 | 2.7 |
| Yoruba | |||||||||||
| Aya-101 | Sent | 31.2 | 1.2 | 8.2 | 5.4 | 3.1 | 57.9 | 2.3 | 23.8 | 13.6 | 6.8 |
| Doc10 | 18.7 | 1.4 | 6.7 | 2.7 | 2.2 | 56.1 | 2.9 | 18.7 | 10.1 | 5.1 | |
| GPT-3.5 | Sent | 21.3 | 1.1 | 7.8 | 5.5 | 3.9 | 52.1 | 2.6 | 12.9 | 5.0 | 3.3 |
| Doc10 | 10.1 | 1.0 | 4.1 | 2.3 | 2.1 | 51.7 | 3.9 | 9.0 | 3.6 | 2.1 | |
| LLaMAX3-SFT1 | Sent | 57.5 | 4.0 | 12.3 | 3.8 | 2.7 | 64.7 | 3.2 | 12.2 | 6.1 | 3.1 |
| Doc10 | 33.8 | 2.7 | 8.1 | 2.4 | 1.8 | 27.2 | 3.2 | 9.1 | 3.4 | 2.0 | |
| LLaMAX3-SFT10 | Doc10 | 52.2 | 4.2 | 10.3 | 2.4 | 1.7 | 71.8 | 4.4 | 10.2 | 5.9 | 2.4 |
| Zulu | |||||||||||
| Aya-101 | Sent | 58.6 | 2.7 | 13.7 | 6.0 | 3.6 | 67.4 | 2.9 | 19.1 | 12.2 | 8.6 |
| Doc10 | 54.9 | 3.1 | 16.1 | 3.9 | 2.7 | 69.0 | 3.3 | 15.7 | 10.7 | 4.5 | |
| GPT-3.5 | Sent | 51.1 | 1.5 | 20.9 | 10.0 | 6.1 | 59.5 | 2.6 | 15.9 | 8.8 | 5.7 |
| Doc10 | 29.2 | 1.3 | 10.1 | 4.0 | 3.2 | 61.1 | 4.0 | 10.8 | 5.3 | 2.9 | |
| LLaMAX3-SFT1 | Sent | 67.5 | 3.3 | 12.4 | 5.5 | 3.6 | 70.5 | 3.6 | 12.5 | 7.2 | 3.8 |
| Doc10 | 33.7 | 2.4 | 8.4 | 2.7 | 2.2 | 31.5 | 3.1 | 8.6 | 3.4 | 2.1 | |
| LLaMAX3-SFT10 | Doc10 | 55.0 | 3.6 | 12.1 | 3.0 | 2.0 | 74.1 | 4.4 | 9.4 | 5.5 | 2.5 |
| Model | Setup | eng X | X eng | ||||||||
| d-CHRF | Fluency | CE | LE | GE | d-CHRF | Fluency | CE | LE | GE | ||
| Amharic | |||||||||||
| Aya-101 | Sent | 37.3 | 3.2 | 9.9 | 3.4 | 3.3 | 65.2 | 2.9 | 19.3 | 11.2 | 5.8 |
| Doc10 | 30.1 | 2.6 | 8.3 | 2.4 | 1.9 | 62.5 | 3.5 | 13.2 | 7.8 | 3.9 | |
| GPT-3.5 | Sent | 22.6 | 1.3 | 8.5 | 4.2 | 3.3 | 47.4 | 2.4 | 8.0 | 3.6 | 2.4 |
| Doc10 | 13.5 | 1.1 | 4.3 | 2.0 | 1.9 | 38.5 | 3.6 | 7.5 | 2.8 | 1.7 | |
| LLaMAX3-SFT1 | Sent | 42.8 | 3.2 | 11.6 | 3.0 | 2.4 | 63.0 | 3.3 | 14.2 | 7.9 | 3.4 |
| Doc10 | 21.7 | 1.7 | 7.3 | 2.2 | 1.8 | 24.2 | 3.0 | 8.6 | 2.9 | 1.9 | |
| LLaMAX3-SFT10 | Doc10 | 37.7 | 2.9 | 8.5 | 2.5 | 1.7 | 65.4 | 4.0 | 10.6 | 5.4 | 2.5 |
| Hausa | |||||||||||
| Aya-101 | Sent | 58.9 | 3.1 | 10.4 | 3.5 | 2.9 | 64.8 | 3.3 | 17.1 | 13.8 | 6.5 |
| Doc10 | 55.0 | 3.8 | 11.7 | 2.5 | 2.1 | 65.5 | 3.5 | 14.0 | 10.9 | 4.5 | |
| GPT-3.5 | Sent | 49.2 | 1.9 | 11.9 | 7.4 | 6.0 | 56.5 | 2.8 | 12.7 | 7.1 | 5.2 |
| Doc10 | 29.7 | 1.6 | 8.8 | 4.5 | 3.7 | 56.3 | 4.1 | 10.3 | 4.9 | 3.9 | |
| LLaMAX3-SFT1 | Sent | 62.4 | 4.0 | 10.1 | 4.1 | 3.0 | 53.5 | 2.9 | 10.3 | 4.7 | 3.4 |
| Doc10 | 29.9 | 2.7 | 7.9 | 2.9 | 2.1 | 27.6 | 2.8 | 8.2 | 3.0 | 2.0 | |
| LLaMAX3-SFT10 | Doc10 | 58.6 | 4.3 | 11.9 | 2.4 | 2.1 | 68.5 | 4.5 | 8.7 | 4.3 | 2.0 |
| Swahili | |||||||||||
| Aya-101 | Sent | 42.4 | 1.5 | 9.0 | 5.0 | 3.3 | 69.1 | 3.2 | 18.4 | 10.9 | 5.4 |
| Doc10 | 51.7 | 2.9 | 9.3 | 2.9 | 1.9 | 68.8 | 3.4 | 15.3 | 10.4 | 4.6 | |
| GPT-3.5 | Sent | 72.6 | 4.7 | 5.0 | 0.8 | 0.5 | 71.5 | 3.4 | 15.3 | 9.4 | 4.6 |
| Doc10 | 72.1 | 4.8 | 5.7 | 0.5 | 0.3 | 73.5 | 4.3 | 11.1 | 7.9 | 4.0 | |
| LLaMAX3-SFT1 | Sent | 67.6 | 3.6 | 12.3 | 4.8 | 3.4 | 67.5 | 3.8 | 12.2 | 6.2 | 3.0 |
| Doc10 | 37.0 | 2.8 | 7.7 | 3.0 | 2.5 | 32.3 | 2.9 | 8.7 | 2.9 | 2.3 | |
| LLaMAX3-SFT10 | Doc10 | 68.3 | 4.2 | 9.5 | 3.2 | 2.1 | 73.1 | 4.4 | 9.1 | 6.1 | 2.6 |
| Yoruba | |||||||||||
| Aya-101 | Sent | 31.4 | 1.9 | 9.0 | 4.8 | 3.7 | 58.5 | 2.8 | 23.2 | 13.5 | 5.4 |
| Doc10 | 22.3 | 2.3 | 7.1 | 2.5 | 2.1 | 55.7 | 3.2 | 13.8 | 8.0 | 4.4 | |
| GPT-3.5 | Sent | 23.0 | 1.8 | 11.0 | 7.6 | 5.0 | 54.0 | 2.7 | 13.4 | 6.6 | 4.5 |
| Doc10 | 12.7 | 1.2 | 6.9 | 2.9 | 2.6 | 53.0 | 3.8 | 9.8 | 4.5 | 2.4 | |
| LLaMAX3-SFT1 | Sent | 55.2 | 4.1 | 9.5 | 3.8 | 2.3 | 57.3 | 3.1 | 9.7 | 4.8 | 3.2 |
| Doc10 | 30.5 | 2.7 | 7.1 | 2.4 | 1.8 | 28.5 | 3.2 | 8.4 | 2.4 | 1.8 | |
| LLaMAX3-SFT10 | Doc10 | 49.3 | 4.1 | 9.0 | 1.5 | 1.2 | 67.7 | 4.4 | 10.6 | 4.4 | 2.1 |
| Zulu | |||||||||||
| Aya-101 | Sent | 58.9 | 2.9 | 12.2 | 4.7 | 3.7 | 67.1 | 3.2 | 17.0 | 13.0 | 5.5 |
| Doc10 | 55.0 | 3.4 | 10.4 | 3.7 | 2.4 | 68.4 | 3.6 | 15.4 | 10.0 | 4.6 | |
| GPT-3.5 | Sent | 53.6 | 1.8 | 22.7 | 7.5 | 5.9 | 59.9 | 2.8 | 16.1 | 11.1 | 7.1 |
| Doc10 | 35.1 | 1.6 | 19.1 | 4.7 | 3.7 | 61.2 | 3.9 | 11.3 | 5.8 | 2.8 | |
| LLaMAX3-SFT1 | Sent | 66.0 | 3.2 | 12.0 | 5.3 | 3.9 | 66.8 | 3.6 | 11.1 | 5.6 | 3.0 |
| Doc10 | 31.7 | 2.4 | 7.7 | 3.1 | 2.4 | 29.8 | 3.2 | 8.0 | 3.1 | 2.2 | |
| LLaMAX3-SFT10 | Doc10 | 60.9 | 3.5 | 11.0 | 4.0 | 2.4 | 71.6 | 4.4 | 8.7 | 5.9 | 3.4 |